OpenStudy (anonymous):

What is the 8th term of the geometric sequence where a1 = 256 and a3 = 16?

OpenStudy (anonymous):

common ratio: r^2 = 16 / 256 = 1 /16 so r = 1/4 so 8th term = a* (1/4)^7 = 256 / (1/4)^7

OpenStudy (chaise):

tn=ar^(n-1) t1=256*r^(1-1) t3=256*r^2 16=256*r^2 1/16=r^2 r=sqrt(1/16) r=1/4 t8=256*(1/4)^7 t8=1/64

OpenStudy (anonymous):

yup